### v1

# Enabling Systematic Generalization in Abstract Spatial Reasoning through Meta-Learning for Compositionality

Abstract
Systematic generalization refers to the capacity to understand and generate novel combinations from known components. Despite recent progress by large language models (LLMs) across various domains, these models often fail to extend their knowledge to novel compositional scenarios, revealing notable limitations in systematic generalization. There has been an ongoing debate about whether neural networks possess the capacity for systematic generalization, with recent studies suggesting that meta-learning approaches designed for compositionality can significantly enhance this ability. However, these insights have largely been confined to linguistic problems, leaving their applicability to other tasks an open question. In this study, we extend the approach of meta-learning for compositionality to the domain of abstract spatial reasoning. To this end, we introduce SYGAR—a dataset designed to evaluate the capacity of models to systematically generalize from known geometric transformations (e.g., translation, rotation) of two-dimensional objects to novel combinations of these transformations (e.g., translation+rotation). Our results show that a transformer-based encoder-decoder model, trained via meta-learning for compositionality, can systematically generalize to previously unseen transformation compositions, significantly outperforming state-of-the-art LLMs, including o3-mini, GPT-4o, and Gemini 2.0 Flash, which fail to exhibit similar systematic behavior. Our findings highlight the effectiveness of meta-learning in promoting systematicity beyond linguistic tasks, suggesting a promising direction toward more robust and generalizable models.
1 Introduction
A fundamental aspect of human cognition is the ability to systematically generalize from known components to novel combinations (Marcus, 2003; Lake et al., 2017). This capacity is particularly evident in language, where an infinite number of new sentences can be constructed and interpreted by extracting meaning from previously acquired expressions and rules (Chomsky, 2002; Szabó, 2012). Similarly, our spatial perception relies on systematic generalization, enabling individuals to compose learned spatial principles into novel configurations (Zhou et al., 2024; Dautriche & Chemla, 2025). For instance, once a person understands how to translate and rotate an object, they can apply these transformations in combination—translating and rotating the object simultaneously—even if they have never encountered such a composed transformation before (Fife et al., 2019).
Despite its central role in human cognition, systematic generalization remains a significant challenge in artificial intelligence (Lake & Baroni, 2018; Loula et al., 2018; Hupkes et al., 2020). While large language models have recently demonstrated notable progress across various domains (OpenAI, 2024; Guo et al., 2025), they often fail to combine acquired knowledge in novel scenarios, demonstrating notable difficulties with systematic generalization (Dziri et al., 2023; Ismayilzada et al., 2025; Gendron et al., 2024). The question of whether neural networks can achieve systematicity has been the subject of extensive debate (Fodor & Pylyshyn, 1988; Brakel & Frank, 2009; Calvo & Symons, 2014, inter alia). Recent research by Lake & Baroni (2023) demonstrates that a transformer-based encoder-decoder model, trained using meta-learning for compositionality (MLC), can achieve human-like systematic generalization in processing instructions expressed in a pseudolanguage. By training the model to combine basic units of pseudolanguage into novel sequences over a stream of dynamically changing grammars, Lake & Baroni (2023) show that this model can effectively generalize to previously unseen compositions of language (see Section 2 for further details). While this approach presents a promising direction for addressing systematicity in neural networks, its applicability beyond linguistic contexts remains an open question.
In this study, we extend the MLC framework proposed by Lake & Baroni (2023) to the domain of abstract spatial reasoning. Drawing inspiration from the abstraction and reasoning corpus (Chollet, 2019), we introduce a new dataset that assesses systematic generalization with respect to basic geometric transformations of two-dimensional objects within grid-based environments (see Figure 1). Using MLC, we train an encoder-decoder model on the task and demonstrate that it generalizes to unseen compositions of geometric transformations, significantly outperforming state-of-the-art LLMs, including GPT-4o (Achiam et al., 2023), o3-mini (OpenAI, 2025), and Gemini 2.0 Flash (DeepMind, 2024), which fail to exhibit comparable systematic behavior. To the best of our knowledge, this is the first application of MLC in abstract spatial reasoning. In summary, our contributions are as follows:
1.
We introduce SYGAR—a new dataset designed to assess systematic generalization in abstract spatial reasoning. This dataset includes basic geometric transformations (e.g., translation, rotation) and their compositions (e.g., translation + rotation) applied to two-dimensional objects in a grid-based environment.
2.
We show that MLC enables models to generalize to unseen compositions of geometric transformations, highlighting its potential beyond linguistic tasks.
3.
We find that models trained via MLC significantly outperform state-of-the-art LLMs such as GPT-4o, o3-mini, and Gemini 2.0 Flash on this task.
2 Background: meta-learning for compositionality
When learning a new language, humans rely on their ability to recombine known words and expressions to interpret novel sentences (Chomsky et al., 1976; De Beule & Bergen, 2006). For instance, someone who understands the meanings of “cats drink water” and “dogs like to play” will typically also understand the meanings of “dogs drink water” and “cats like to play” (Hinzen et al., 2012). Whether language models possess a comparable degree of systematicity remains an open question, as current models, including large language models, still struggle with tests of systematic generalization (Ismayilzada et al., 2025; Dziri et al., 2023). To address these limitations, Lake & Baroni (2023) propose meta-learning for compositionality (MLC), a framework designed to model human-like systematic generalization in learning pseudolanguage instructions. Through a series of experiments, the authors show that models trained via MLC can achieve levels of systematicity comparable to those of humans when interpreting previously unseen pseudolanguage inputs.
Task setup.
In their study, Lake & Baroni (2023) examine few-shot compositional tasks in which instructions, represented as sequences of pseudowords (e.g., “dax,” “lug,” “fep”), must be mapped to corresponding sequences of abstract symbols (see Figure 2 for an example). To understand the meaning of such instructions, an interpretation grammar needs to be deduced from a limited number of study examples. This grammar maps pseudowords to their symbolic representation through a set of compositional rewrite rules. For instance, if “dax” corresponds to a green circle, “dax fep” to three green circles, and “zup” to a red circle, then “zup fep” would denote three red circles. Importantly, the examples are designed to be highly systematic, progressing from primitive mappings to more complex compositions. The core challenge lies in the ability to generalize systematically, i.e., to reuse and combine components from the study examples (left side of Figure 2) to generate correct outputs for novel query instructions (right side of Figure 2).
Algorithmic approach.
To achieve systematic generalization in the instruction-learning task, Lake & Baroni (2023) train a transformer-based encoder-decoder model through meta-learning for compositionality. The key idea is to train the model on a dataset of dynamically changing interpretation grammars, where the mappings from input sequences to output symbols differ across training samples. This forces the model to rely on the information conveyed in the study examples to infer the appropriate grammar of a given sample, rather than memorizing static input-output mappings across the dataset. This flexibility enables the model to adjust to novel scenarios governed by new sets of examples and rules. Moreover, the compositional structure of both study examples and queries encourages the model to internalize mechanisms for composing elements presented in the study examples. After training the model over a set of 100,000 distinct interpretation grammars, it demonstrates the capacity to generalize to previously unseen instructions and grammars. For specific details regarding training procedures, we refer to the original paper (Lake & Baroni, 2023).
While Lake & Baroni (2023) also evaluate MLC on COGS (Kim & Linzen, 2020) and SCAN (Lake & Baroni, 2018), which test systematic lexical generalization to novel word combinations, their experiments are confined to the linguistic domain. In the following section, we show how MLC can be extended to support systematic generalization in abstract spatial reasoning, demonstrating its potential beyond linguistic tasks.
3 Method
3.1 SYGAR: a dataset for SYstematic Generalization in Abstract spatial Reasoning
To test systematicity in abstract spatial reasoning, we leverage the closure property of combined geometric transformations, where the composition of two valid transformations—such as translation, rotation, and reflection—yields another valid geometric transformation (Brannan et al., 2011). Drawing inspiration from the abstraction and reasoning corpus (Chollet, 2019), we design a task in which abstract objects, defined in a two-dimensional grid environment, are subjected to basic geometric transformations and their compositions (see Figure 1 for examples). We use fixed-size grids, each of which can be represented as a two-dimensional array of integers, where different values correspond to distinct colors. We use integers from 0 to 9, with 0 denoting a black background and the remaining integers mapping to unique colors (see Appendix A.1 for more details). Objects are defined based on color connectivity; that is, each object comprises a group of connected cells sharing the same color. Connectivity is determined by the Moore neighborhood (Bays, 2010), meaning that cells are considered connected if they are directly or diagonally adjacent. Each grid contains either one or two objects. A transformation is represented as a pair of grids, with the input grid displaying the objects before, and the output grid showing them after the geometric transformation. Each transformation affects only one of the objects in the grid. For example, in Figure 1(a), a single L-shaped yellow object is translated one step downward. In Figure 1c, a square blue object in the bottom-right expands toward the neighboring top row. Objects never occlude one another nor extend beyond the boundaries of the grids.
We limit our dataset to five basic geometric transformations and their compositions: i) translations, ii) rotations, iii) reflections, iv) extensions, and v) color changes. For our experiments, we further constrain the configurations of these transformations to establish a controlled setup. Translations are limited to movements of one cell to the right or one cell downward. Rotations are restricted to 90 degrees clockwise or counterclockwise around the top-left corner of the object. We consider horizontal and vertical reflections across the object’s central axis. Extensions mean that the object grows in a certain direction, and are limited to neighboring cells either leftward or upward. Color changes are restricted to changing the object’s color to either red or orange. For detailed definitions of each transformation, please refer to Appendix A.2.
To signal which objects undergo which transformations, we consider three types of indicators: i) shape-based transformations, which affect objects of a particular shape; ii) color-based transformations, which affect all objects of a specific color; and iii) neighbor-based transformations, where objects are transformed when a second, indicator object is present. For instance, in Figure 1, all L-shaped objects (similar to the object in Figure 1(a)) undergo a one-step downward translation. All green objects undergo a horizontal reflection, and any object sharing a grid with the gray diagonal object (e.g., as seen in Figure 1c) expands into the neighboring top row. This indicator-based approach enables the definition of transformation compositions. For example, objects that are both L-shaped and green undergo a one-step downward translation together with a horizontal reflection (see Figure 1d for an example). We also define different levels of composition: level 1 combines two indicators (e.g., when an object matches the indicated shape and color, but lacks a the proximity to a neighboring object, as illustrated in Figure 1d), while level 2 combines all three indicators, specifying the object’s shape, color, and proximity to an indicator object (see Figure 1g).
To test systematicity, we present examples of primitive transformations and their compositions, and evaluate models on previously unseen combinations of indicators. For instance, in Figure 3, models are asked to infer the correct transformation for a previously unseen level-2 composition of indicators, given study examples of primitive transformations and their level-1 compositions. Conceptually, our setup is similar to the few-shot compositional task introduced by Lake & Baroni (2023) (see Section 2), but it replaces the lexical interpretation grammar with a visual interpretation grammar. Specifically, models need to infer which indicator maps to which transformation, and how to compose them to deduce the correct final transformation. For a detailed description of how we algorithmically generate dataset samples, please refer to Appendix A.3.
3.2 Meta-Learning for compositionality in abstract spatial reasoning
To systematically generalize from known geometric transformations to previously unseen transformation compositions, we extend the meta-learning for compositionality (Lake & Baroni, 2023) framework described in Section 2. As in the original MLC approach, we train a transformer-based encoder-decoder model on a dataset of dynamically changing interpretation grammars. However, instead of mapping pseudolinguistic instructions to sequences of abstract symbols, we consider a visual interpretation grammar that associates visual indicators (object shape, color, or proximity to an indicator object) with specific geometric transformations, as described in Section 3.1. An episode is defined as a set of study examples that demonstrate the underlying grammar, along with query inputs for which the correct outputs must be inferred (see Figure 3 for an illustration). By training over a series of episodes with changing visual interpretation grammars, the model needs to abstract and recombine information from the study examples in order to predict the correct output for previously unseen query grids.
Encoding and positional embedding.
Each episode is presented to the model as a sequence of input-output grid pairs (study examples), followed by a query input grid, for which the model must generate the corresponding output grid (see Figure 3). To encode the two-dimensional grids, we divide each grid into patches (left to right, top to bottom), yielding 25 patches per grid (Dosovitskiy et al., 2021). Each patch is mapped to a unique embedding vector. Since each grid cell can take integer values from 0 to 9, a patch can yield up to 10,000 distinct configurations, resulting in 10,000 possible embedding vectors. Two special tokens, and , are introduced to mark the boundaries between study examples and the input-output grids, respectively. The decoder vocabulary comprises two additional tokens for the start and end of a sequence (SOS and EOS). To encode positional information, we use standard learnable 1D positional embeddings that capture the order of grid pairs, as well as a second set of learnable 2D positional embeddings applied to grid patches. These 2D embeddings are decomposed into separate row and column components, which are added to each patch embedding to capture two-dimensional spatial information.
Training procedure.
The model is trained on a large set of episodes, each defined by a unique visual interpretation grammar. In each episode, the model is provided with a sequence of study examples and tasked with predicting the output grid for a given input query (see Figure 3). Following Lake & Baroni (2023), we include an auxiliary copy task during training, in which the model must also reproduce the output grids of each study example. We employ a model with three layers each in the encoder and decoder, eight attention heads per layer, input and hidden embeddings of size 128, a feedforward hidden size of 768, and GELU (Hendrycks & Gimpel, 2016) activations. In total, the model has 5.7 million trainable parameters. To promote robustness in the decoder, we introduce minor perturbations by randomly altering the color of individual cells in the target output query with a small probability (0.001). Unlike Lake & Baroni (2023), we do not incorporate systematic noise to model inductive biases observed in human learning. Further implementation details regarding the training procedure and hyperparameters can be found in Appendix B.
4 Experimental setup
4.1 Task setup
We consider two different task setups in this work. The first is a standard few-shot learning scenario, in which models are required to generate an output grid for a given query input that undergoes a level-2 transformation composition. This prediction is based on three examples that demonstrate the same level-2 transformation. A visual representation of this setup is provided in Figure 4 in the Appendix. This task evaluates the model’s ability to infer geometric transformations from a limited set of illustrative examples.
The second setup focuses on systematicity and differs from the first in the type of few-shot examples presented. As mentioned in Section 3.1, the idea is to test whether models can infer novel compositions from known geometric transformations. For this, we replace the level-2 few-shot examples with a set of primitive transformations plus level-1 transformation compositions, and ask the model to predict the previously unseen level-2 transformation composition, as illustrated in Figure 3. Specifically, we present six primitive transformations—two examples for each indicator (shape-based, color-based, neighbor-based)—and six level-1 transformation compositions, two examples for each first-level indicator composition (shape+color, shape+neighbor, color+neighbor).
We generate 100,000 episodes, each comprising three few-shot examples for the standard few-shot learning task, 12 systematic study examples for the systematicity setup, and ten query input-output grid pairs demonstrating the final level-2 transformation composition. Each episode is characterized by a unique visual interpretation grammar. For instance, in one episode, yellow objects are translated downward by a single cell, while in another, yellow objects are reflected horizontally. This challenges models to generalize to novel scenarios by inferring the correct transformations based on the provided examples. To train our encoder-decoder model via MLC, we split the data into 82,908 training, 8,546 validation and 8,546 test episodes. Importantly, the data splits are constructed such that the geometric transformations involved in the final query level-2 compositions differ between the training and evaluation sets. For instance, while the model is trained on all basic transformations and a series of transformation compositions (e.g., translation+rotation+reflection), it is tested on compositions not seen during training (e.g., translation+rotation+extension). For comprehensive statistics of the dataset splits, please refer to Appendix A.4.
4.2 Language models
In addition to the model trained via MLC, we evaluate three state-of-the-art LLMs on the test set of our proposed dataset: GPT-4o (Achiam et al., 2023), o3-mini (OpenAI, 2025), and Gemini 2.0 Flash (DeepMind, 2024). To textually prompt the models for a given episode, we represent grids as two-dimensional arrays, consistent with prior work (Moskvichev et al., 2023). We also test a multimodal setup in which both an image of the study examples and the input query are provided alongside the text prompt (text+image). Due to financial constraints, each model is evaluated on a single test query for each episode across the 8,546 episodes in the test set. All textual and visual prompts, specific model versions, and decoding parameters are detailed in Appendix C.2.
4.3 Evaluation metrics
To assess the quality of the generated output grids, we employ three different evaluation metrics: i) accuracy, ii) color accuracy, and iii) shape accuracy. For accuracy, we use exact match, i.e., an output is considered accurate if all cells in the predicted output grid correspond to those in the target grid. Color accuracy measures whether the predicted objects in the output grid match the colors of the objects in the target grid, regardless of their shape or location. Shape accuracy, on the other hand, evaluates whether the predicted objects share the same shape as those in the target grid, irrespective of color or location. For formal definitions of these metrics, please refer to Appendix C.1.
| Model | Accuracy [%] | Color Acc. [%] | Shape Acc. [%] | |
| 3-Shot | GPT-4o (text-only) | |||
| GPT-4o (text+image) | ||||
| Gemini 2.0 Flash (text-only) | ||||
| Gemini 2.0 Flash (text+image) | ||||
| o3-mini | ||||
| MLC | 99.92 | 100.00 | 99.92 | |
| Systematicity | GPT-4o (text-only) | |||
| GPT-4o (text+image) | ||||
| Gemini 2.0 Flash (text-only) | 99.68 | |||
| Gemini 2.0 Flash (text+image) | ||||
| o3-mini | ||||
| MLC | 78.26 | 80.49 |
5 Results
In Table 1, we report the performance of our model trained via MLC, alongside the LLMs we evaluate on the two task setups, as described in Section 4.1.
Standard few-shot learning task.
We begin by examining model performance on the three-shot learning task. In this setup, models are provided with three input-output examples that illustrate the final transformation (see Figure 4 in the Appendix for a visual illustration). Despite this guidance, LLMs such as GPT-4o and Gemini 2.0 Flash exhibit notable difficulties with this task. GPT-4o achieves a maximum accuracy of 22.28% (text-only), while Gemini 2.0 Flash performs marginally better with an accuracy of up to 30.08% (text-only). Although the long-chain-of-thought model o3-mini achieves a higher performance of up to 64.04%—performing best among all LLMs—its accuracy remains modest given the simplicity of the transformations involved. In contrast, our encoder-decoder model trained via MLC achieves an accuracy of up to 99.92%, substantially surpassing all other models. When comparing the models’ overall accuracy with their shape and color accuracy, we find that all models perform nearly perfectly in predicting object color. Notably, for GPT-4o and Gemini 2.0 Flash, shape accuracy is significantly higher than exact match accuracy. This discrepancy suggests that while these models are often able to predict the correct shape of an object, they frequently fail to accurately predict its final position. Interestingly, both GPT-4o and Gemini 2.0 Flash show a decline in accuracy when an additional visual input is included alongside the textual prompt. We hypothesize that this performance degradation stems from the added complexity introduced by the image, either due to modality alignment challenges (Masry et al., 2025) or limitations in leveraging the visual content for reasoning.
Systematicity task.
In the systematicity task, models are asked to infer the correct final transformation composition from a set of study examples that represent more basic, decomposed transformations (see Figure 3 for an example). As shown in Table 1, all LLMs perform poorly on this task. For instance, GPT-4o achieves a maximum accuracy of 0.99% (text-only), while Gemini 2.0 Flash reaches up to 2.66% (text-only). Interestingly, o3-mini, the model that performed best among all LLMs on the standard few-shot learning task, performs worst in this setting, with an accuracy of only 0.53%. In contrast, our encoder-decoder model trained via MLC achieves an accuracy of up to 78.26%, significantly outperforming all other LLMs despite having orders of magnitude fewer parameters. Importantly, as described in Section 4.1, this model has never seen the specific level-2 compositions of geometric transformations in the test queries during training but was instead optimized on a distinct set of transformation compositions (see Table 5 in the Appendix). Consistent with our findings from the 3-shot learning task, models generally succeed in predicting the correct object colors. However, shape accuracy declines considerably in this setting, indicating that models struggle to predict the correct shape of the output object after the transformation. A qualitative example of the models’ predictions can be found in Figure 3, with additional examples provided in Appendix D. The strong performance of our model trained via MLC highlights the effectiveness of this training strategy in promoting systematic generalization to novel transformation compositions. The model not only learns to infer a visual interpretation grammar from a limited number of study examples but also generalizes to novel transformation compositions that it has never encountered during training.
5.1 Consistency across data splits
To ensure that the strong performance of MLC, as reported in Table 1, is not the result of a favorable data split, we train and evaluate the model on three additional, independently generated data splits for each task configuration—resulting in four distinct models per task setup. Detailed descriptions of these data splits are provided in Table 5 in the Appendix. Table 2 summarizes the average accuracy and corresponding standard deviation across all four splits. For the standard three-shot learning task, MLC consistently achieves high accuracy, with a mean of 98.78% and a standard deviation of 1.99%. Similarly, in the systematicity task, the model demonstrates robust generalization, achieving an even higher average accuracy than on the initial data split, with a mean of 86.73% 6.03%.
| Model | Accuracy [%] | Color Acc. [%] | Shape Acc. [%] |
| MLC (3-Shot) | |||
| MLC (Systematicity) | 86.73 6.03 | 99.36 0.70 | 87.55 5.45 |
| - no copy task | |||
| - no primitives | 99.56 0.50 | ||
| - no level-1 compositions |
Ablation studies.
To gain deeper insights into the factors influencing model performance, we conduct a series of ablation studies. First, we evaluate the impact of removing the auxiliary copy task from the training objective—a setup in which the model is trained not only to predict the output grid for a given input query but also to reproduce the output grid of each study example (refer to Section 3.2). Removing this auxiliary task results in a notable decrease in accuracy from 86.73% 6.03% to 69.05% 9.23%. This decline underscores the importance of the copy task in promoting systematic generalization, aligning with the findings of Lake & Baroni (2023). Subsequently, we assess the role of study examples in model performance. Removing primitive transformations from the study examples (see Figure 3) results in a moderate reduction in performance, with an average accuracy of 75.27% 12.95%. This suggests that examples involving only level-1 transformation compositions are, to some extent, sufficient for allowing the model to generalize to more complex level-2 compositions. However, removing level-1 transformation compositions leads to a severe performance degradation, reducing accuracy to 21.01% 19.07%. We hypothesize that this is due to the increased complexity of composing three primitive operations directly into a level-2 transformation, as opposed to building on intermediate level-1 compositions.
In conclusion, our experiments highlight the potential of MLC beyond linguistic tasks. The presented results demonstrate that MLC enables systematic generalization to novel transformation compositions, significantly outperforming current state-of-the-art LLMs.
6 Related work
Meta-Learning.
Meta-learning aims to improve a model’s ability to adapt to novel tasks by leveraging experience over multiple training episodes (Thrun & Pratt, 1998; Hospedales et al., 2022). It has been successfully applied to various tasks, such as few-shot learning (Mishra et al., 2018), continual learning (Javed & White, 2019; Lee et al., 2023; Irie et al., 2025), and reinforcement learning (Duan et al., 2016; Wang et al., 2017; Mishra et al., 2018). Related to our work, meta-learning has been used to improve systematic generalization. Lake & Baroni (2018) showed that traditional sequence-to-sequence models struggle with compositional skills, but incorporating meta-learning can significantly improve performance (Lake, 2019; Conklin et al., 2021). Recent work argues that giving models the opportunity to practice skills via meta-learning is crucial for addressing challenges such as systematic generalization, among others (Irie & Lake, 2024). Our method builds on meta-learning strategies inspired by Lake & Baroni (2023), extending them to the domain of abstract spatial reasoning.
ARC-like puzzles.
The abstraction and reasoning corpus (ARC) (Chollet, 2019) is a benchmark designed to evaluate a model’s capacity to generalize to novel scenarios with limited to no prior knowledge. Based on a set of few-shot examples, models are required to infer transformations of abstract objects or patterns within two-dimensional grids. Unlike ARC, which encompasses a broad range of complex transformations, our work deliberately narrows the scope to the five fundamental geometric transformations described in Section 3.1, focusing instead on the aspect of systematicity. Several ARC variants have been proposed, including 1D-ARC (Xu et al., 2023), Mini-ARC (Kim et al., 2022), ConceptARC (Moskvichev et al., 2023) and MC-LARC (Shin et al., 2024). However, our dataset, SYGAR, is the first to focus on the aspect of systematicity in abstract spatial reasoning.
7 Conclusion
In this work, we extend the meta-learning for compositionality framework proposed by Lake & Baroni (2023) to the domain of abstract spatial reasoning. To this end, we introduce SYGAR—a novel dataset designed to evaluate systematicity in this field. Our experiments demonstrate that models trained via MLC can systematically generalize to novel compositions of geometric transformations. Moreover, MLC consistently outperforms current state-of-the-art LLMs, which do not exhibit comparable systematic behavior. Our findings suggest that MLC presents a promising direction for enabling systematic generalization in language models across diverse domains.
Reproducibility statement
To ensure the reproducibility of our work, we make all code publicly available at: https://github.com/mainlp/SYGAR. This enables users to reproduce the data described in Section 3.1 and train models via MLC for the task, as outlined in Section 3.2. Details about the training procedures and hyperparameters are provided in Section 3.2 and Appendix B. Additionally, we include an exemplary subset of input queries and corresponding API responses from the evaluated LLMs as part of the supplementary material. Specifics on prompts, model versions, and decoding parameters are given in Appendix C.2. Further details about the datasets can be found in Section 3.1, Section 4.1, and Appendix A. Finally, Appendix B.2 outlines the software and computational resources used for model training.
Acknowledgments
We express our gratitude to the members of the MaiNLP lab for their invaluable feedback. Furthermore, we thank the anonymous reviewers for their insightful comments and suggestions. We gratefully acknowledge that experiments involving API calls to GPT-4o and o3-mini were supported by a compute grant from OpenAI. The authors also acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b217dd. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. Finally, we acknowledge the support for BP through the ERC Consolidator Grant DIALECT 101043235.
Appendix A Dataset
In this work, we present SYGAR, a dataset designed to study systematicity in abstract spatial reasoning. As outlined in Section 3.1, SYGAR evaluates a model’s capacity to systematically generalize learned geometric transformations (e.g., translation, rotation) of two-dimensional objects to novel compositions of these transformations (e.g., translation+rotation). The subsequent sections offer a detailed description of the dataset, including formal definitions of the grid-based environment and the set of transformations it includes.
A.1 Grid setup
We define the structure of the grid environment and the notion of objects within it. Each grid is represented as a matrix , where each element corresponds to a cell with a discrete color value. Objects are defined based on color connectivity using the Moore neighborhood (Bays, 2010).
[Grid & Object] Let be a matrix, referred to as a grid, where each element . The value represents a background cell, and values represent object colors.
An object is a set of coordinates
such that each satisfies , and the elements in form a single connected component.
Two elements and are considered connected if:
The object size is the number of elements it contains.
We define the following color mapping: , , , , , , , , , and .
A.2 Geometric transformations
We formally define the five basic geometric transformations used in our dataset: translation, rotation, reflection, extension, and color change. Each transformation operates on objects within the grid environment as defined in Appendix A.1. A transformation is considered valid if all transformed coordinates lie within the grid bounds and do not overlap with existing objects in the original grid.
Translation. Moves an object by one cell along a specified direction (downward or rightward). A formal definition is given in the text box below.
[Translation] Let be an object in a grid , and let be the translation direction (downward or rightward).
The translated object is:
The translation is valid if:
Rotation. Rotates an object clockwise or counterclockwise around the top-left of its bounding box. A more formal definition is given in the text box below.
[Rotation] Let be an object in a grid , and let be the top-left coordinate of the object’s bounding box. Let be the rotation angle.
For each , let the relative offset be:
Let the rotated offset be given by:
Then the rotated object is:
The rotation is valid if:
Reflection. Reflects an object across its vertical or horizontal axis, reversing the relative positions of its coordinates while preserving overall structure.
[Reflection] Let be an object in a grid , and let indicate the axis of reflection.
Let:
Then the reflected object is:
The reflection is valid if:
Extension. Adds a new cell in the upward or leftward direction for each coordinate in the object.
[Extension] Let be an object in a grid , with color . Let indicate the extension direction.
Let the set of new cells adjacent to the object in direction be:
Then the extended object is:
The extension is valid if:
All new cells are assigned the color of the original object:
Color change. Alters the color of an object to either red or orange, without changing its structure or position.
[Color Change] Let be an object in a grid , with color . Let be the new color (representing red or orange).
The resulting grid is given by:
A.3 Dataset generation
To generate episodes that comprise primitive transformations, level-1 transformation compositions, and level-2 transformation compositions, we developed a script that systematically generates the corresponding input-output grid pairs for each transformation. The complete code repository for data generation is publicly available at: https://github.com/mainlp/SYGAR. In the following, we provide a brief overview of the procedure used to generate input-output grid pairs for each sample within an episode. As detailed in Section 3.1 and Appendix A.2, we consider five basic geometric transformations, along with three types of transformation indicators: shape-based, color-based, and neighbor-based. These allow us to define a total of ten distinct transformation triplets, each mapping the indicators to corresponding transformations (e.g., shape-based: translation, color-based: reflection, neighbor-based: extension). For each episode, a transformation triplet is uniformly sampled from this set to define the visual interpretation grammar of the episode. Once the transformations are determined, we randomly assign a specific shape for the shape-based transformation, a specific color for the color-based transformation, and an indicator object for the neighbor-based transformation. Importantly, the indicator object is constrained to neither share the shape associated with the shape-based transformation nor the color linked to the color-based transformation.
Using these specifications, we generate input-output grid pairs representing primitive, level-1, and level-2 transformations. For each transformation mapping, we randomly place an object on a grid, ensuring it possesses the designated shape, color, and/or proximity to the indicator object as required. The specified transformation is then applied to this object. If the resulting transformed object remains within the grid bounds and does not overlap with any other object, the corresponding input-output grid pair is accepted as a valid sample for the episode. Otherwise, a new object location is sampled and the process is repeated until a valid pair is obtained. Finally, we make sure that each episode follows a unique grammar, i.e., that no two combinations of shape, color, and indicator objects correspond to the same set of transformations within the dataset.
A.4 Dataset statistics
Table 5 presents detailed statistics for the datasets used in this study. As outlined in Section 5.1, we train and evaluate models via MLC across four distinct dataset splits to mitigate the influence of randomness in the data split process. The table includes the number of training, validation, and test samples for each split. Additionally, it provides information on the query transformation compositions present in the training and test sets, along with the frequency of each basic geometric transformation within the train dataset.
Appendix B Training details
As outlined in Section 3.2, we use a transformer-based encoder-decoder model trained using MLC to predict the correct output grid for a given input query, given a set of study examples. Specifically, we generate a dataset of 100,000 episodes and split it into train, validation and test sets (for more information see Section 4.1 and Table 5). The model is optimized using cross-entropy loss, averaged over the predicted patch embeddings, as described in Section 3.2. To place greater emphasis on non-background regions, patches corresponding exclusively to black cells are down-weighted by a factor of during loss computation.
Each episode includes a collection of study examples and queries. In the standard few-shot learning task (Section 4.1), the model receives three input-output grid pairs, along with the input query. For the systematicity task, 12 systematic study examples are provided. In both tasks, the model is required to predict the correct output grid for ten distinct input queries.
Training is conducted over 200 epochs with a batch size of 200 for the standard few-shot learning task (i.e., queries per batch), and over 300 epochs with the same batch size for the systematicity task. A learning rate of 0.01 is used in both cases. Following the approach of Lake & Baroni (2023), we apply a warm-up phase during the first episode, beginning with a learning rate of , followed by a linear decay to over the course of training. Additional hyperparameter settings are provided in Section B.1 and summarized in Table 3.
| Parameter | Value |
| number layers in decoder | 3 |
| number layers in decoder | 3 |
| number of attention heads | 8 |
| hidden dimension | 128 |
| feedforward hidden size | 768 |
| learning rate | 0.01 |
| learning rate after training | |
| dropout | 0.0 |
| weight decay | 0.01 |
| noise probability | 0.001 |
| gradient accumulation over batches | 2 |
| background patch loss weight | 0.2 |
B.1 Hyperparameters
To identify suitable hyperparameters for model training, we conduct Bayesian search over a predefined range of values: learning rate , final learning rate after linear decay , dropout rate , gradient accumulation over batches, cell color perturbation probability , feedforward hidden dimension , loss weighting for background (all-black) patches , number of encoder layers , and number of decoder layers .
For the hyperparamter search, the model is trained for 40 epochs on the systematicity task and evaluated on its corresponding validation set. Across 25 independent runs, we select the configuration that achieves the highest validation accuracy. The final hyperparameter settings, presented in Table 3, are employed consistently across both task setups.
B.2 Implementation details
All experiments were conducted using PyTorch (Paszke et al., 2019) as the primary development framework. Comprehensive details regarding supporting software and versioning are available in our code repository. Experiments were executed on NVIDIA A100 and H200 GPUs. Training models with MLC on the standard three-shot learning task over 200 epochs required approximately 40 GPU hours on a single A100 GPU. For the systematicity experiments with 12 study examples, training over 300 epochs on the designated dataset consumed roughly 100 GPU hours on a single H200 GPU.
Appendix C Experiment details
This section provide further details regarding our experimental setup. Specifically, Section C.1 presents formal definitions of the evaluation metrics used to assess the performance of the models studied in this work, while Section C.2 outlines additional information on how we interact with API-based LLMs.
C.1 Evaluation metrics
As described in Section 4.3, we use three different evaluation metrics to assess model performance in this study: i) exact match accuracy, ii) color accuracy, and iii) shape accuracy. These metrics are formally defined based on the grid-based environment and the concept of an object , as specified in Definition A.1.
Let denote the target and predicted grids, respectively. Each cell (or ) contains an integer in , where 0 represents the background and values from 1 to 9 correspond to cells occupied by colored objects. The set of objects—defined as maximal connected cells of a consistent color under the Moore neighborhood (see Section 3.1)—extracted from and are denoted and , respectively. For each object in grid , we assign a color label and define its normalized shape as follows:
where
This transformation “anchors” the object to the top-left corner by translating it to a coordinate system with its minimum row and column indices set to zero.
Accuracy.
The exact match accuracy evaluates whether the predicted grid is identical to the target grid on a cell-by-cell basis:
In other words, this metric yields a value of 1 if and only if the entire predicted grid matches the target grid exactly, i.e., . The mean accuracy over the dataset is then defined as:
Color accuracy.
Color accuracy assesses whether the predicted grid contains the same number of objects of each color as the target grid, irrespective of their locations or shapes. For a given color , let
denote the number of objects of color in grid . Then, color accuracy is defined as:
where is the indicator function, returning 1 if the condition is satisfied for all colors and 0 otherwise. The mean color accuracy over the dataset is given by:
Shape accuracy.
Shape accuracy measures the agreement in object shapes between the predicted and target grids, independent of color and position. For each object in a grid , we consider its normalized shape as defined in Equation 1. The count of objects with a specific normalized shape in grid is given by:
Accordingly, shape accuracy is defined as:
That is, the predicted grid has perfect shape accuracy if the number of objects corresponding to each normalized shape is identical to that in the target grid . Finally, the mean shape accuracy over the dataset is given by:
C.2 Model information
As described in Section 4.2, we evaluate three different LLMs in addition to our model trained via MLC. Specifically, we assess the performance of GPT-4o (Achiam et al., 2023) (version gpt-4o-2024-08-06111https://platform.openai.com/docs/models/gpt-4o), o3-mini (OpenAI, 2025) (version o3-mini-2025-01-31222https://platform.openai.com/docs/models/o3-mini), and Gemini 2.0 Flash (DeepMind, 2024) (version gemini-2.0-flash-001333https://ai.google.dev/gemini-api/docs/models#gemini-2.0-flash). All models are accessed via their respective batch APIs, enabling us to process multiple samples per request. Unless otherwise specified, we employ the default API settings. For GPT-4o and o3-mini, this corresponds to a temperature and top_p value of .444https://platform.openai.com/docs/api-reference/chat/create Due to financial constraints, the o3-mini model is configured with a “low” reasoning effort. For Gemini 2.0 Flash, the provider does not disclose default parameter settings.
Prompts.
The complete set of prompts used in our evaluations is presented in Figures 11 through 14. To ensure consistency and facilitate meaningful comparisons, we apply the same prompts across all models. The standard few-shot learning prompt appears in Figure 11, while the prompt used for the systematicity task is shown in Figure 13. For Gemini 2.0 Flash, we add the instruction: “Do not generate any code to solve the task” to the output requirements, as the model otherwise does not adhere to the required output format. As outlined in Section 4.2, we additionally evaluate GPT-4o and Gemini 2.0 Flash in a multimodal configuration, in which both an image of the study examples and the input query are provided alongside the text prompt (text+image). The multimodal prompt for the few-shot learning task is shown in Figure 12, with the accompanying image illustrated in Figure 9. The corresponding multimodal prompt for the systematicity task is depicted in Figure 14, with the associated image presented in Figure 10. For the textual prompts, we represent grids as two-dimensional arrays, consistent with prior work (Moskvichev et al., 2023)). For instance, the final query input grid in Figure 4 would be represented as:
Model responses are parsed using regular expressions to identify the expression “output:“, followed by a two-dimensional array of the form “[[]]”, as specified in the input prompt. If a response does not contain this pattern, it is excluded from further analysis and omitted from accuracy computations. Table 4 summarizes the proportion of valid responses for each model.
Appendix D Additional results
In this section, we present additional results for the experiments conducted in this study. First, we present additional qualitative results related to the model predictions on the standard few-shot learning and the systematicity task. Figures 4 through 6 illustrate representative episodes from the standard few-shot learning task. Model predictions are shown adjacent to each query, with results for GPT-4o and Gemini 2.0 Flash corresponding to text-only prompts. Across all three episodes, the model trained using MLC consistently predicts the correct output grid. In contrast, GPT-4o and Gemini 2.0 Flash frequently fail to identify the correct transformation—either misrepresenting the shape of the transformed object or incorrectly predicting its final position. Notably, o3-mini successfully predicts the correct output for the episodes in Figures 5 and 6, but fails on the example in Figure 4. Figures 7 and 8 highlight episodes from the systematicity task. As shown, all baseline models fail to produce accurate transformations, often misplacing the transformed object within the grid. In contrast, the model trained via MLC consistently predicts the correct transformation.
| Model | Valid Responses (3-Shot) | Valid Responses (Systematicity) |
| GPT-4o (text-only) | ||
| GPT-4o (text+image) | ||
| Gemini 2.0 Flash (text-only) | ||
| Gemini 2.0 Flash (text+image) | ||
| o3-mini | ||
| MLC |
Response rates.
As outlined in Section C.2, LLMs we evaluate are instructed to present their final output grid predictions using the keyword “output:”, followed by a two-dimensional array of size in the format “[[]]”. Responses that do not conform to this expected pattern are excluded from subsequent analyses and are not included in accuracy calculations. Table 4 provides an overview of the proportion of valid responses for each model. In the standard few-shot learning setting, all models demonstrate very high valid response rates, exceeding 99%. However, in the systematicity task, a slight decrease in valid responses is observed for Gemini 2.0 Flash when additional visual input (text+image) is introduced, with the rate falling to 94.09%. More significantly, GPT-4o exhibits a notable drop in valid response rate to 77.24% under multimodal conditions. We hypothesize that this decline may be attributed to the increased context length resulting from the additional image input.
| Data Split | No. Episodes | Query Transformations | Basic Transformations | |||
| Set | No. | Type | Composition | Transformation | Freq. | |
| seed 1860 | Train | 82908 | Train | translation+reflection+coloring | red coloring | 35828 |
| Val | 8546 | reflection+rotation+extension | orange coloring | 35819 | ||
| Test | 8546 | translation+reflection+rotation | down translation | 23398 | ||
| translation+rotation+coloring | right translation | 27021 | ||||
| reflection+coloring+extension | leftward extension | 22140 | ||||
| reflection+rotation+coloring | upward extension | 21806 | ||||
| translation+coloring+extension | cw. rotation | 19551 | ||||
| rotation+coloring+extension | ccw. rotation | 19394 | ||||
| Test | translation+rotation+extension | horizontal reflection | 21967 | |||
| translation+reflection+extension | vertical reflection | 21800 | ||||
| seed 1870 | Train | 83481 | Train | translation+rotation+extension | red coloring | 27603 |
| Val | 8259 | translation+reflection+rotation | orange coloring | 27525 | ||
| Test | 8260 | reflection+rotation+extension | down translation | 31385 | ||
| reflection+coloring+extension | right translation | 36126 | ||||
| translation+reflection+extension | leftward extension | 26501 | ||||
| translation+rotation+coloring | upward extension | 25913 | ||||
| translation+reflection+coloring | cw. rotation | 15421 | ||||
| translation+coloring+extension | ccw. rotation | 15283 | ||||
| Test | rotation+coloring+extension | horizontal reflection | 22366 | |||
| reflection+rotation+coloring | vertical reflection | 22320 | ||||
| seed 1880 | Train | 80035 | Train | translation+coloring+extension | red coloring | 25850 |
| Val | 9982 | translation+rotation+extension | orange coloring | 25832 | ||
| Test | 9983 | translation+rotation+coloring | down translation | 31385 | ||
| reflection+rotation+extension | right translation | 36126 | ||||
| translation+reflection+coloring | leftward extension | 24821 | ||||
| translation+reflection+extension | upward extension | 24147 | ||||
| translation+reflection+rotation | cw. rotation | 19734 | ||||
| rotation+coloring+extension | ccw. rotation | 19594 | ||||
| Test | reflection+rotation+coloring | horizontal reflection | 16331 | |||
| reflection+coloring+extension | vertical reflection | 16285 | ||||
| seed 1890 | Train | 80557 | Train | translation+coloring+extension | red coloring | 30227 |
| Val | 9721 | translation+reflection+rotation | orange coloring | 30255 | ||
| Test | 9722 | rotation+coloring+extension | down translation | 23279 | ||
| translation+reflection+coloring | right translation | 24789 | ||||
| reflection+rotation+extension | leftward extension | 26483 | ||||
| translation+reflection+extension | upward extension | 26277 | ||||
| reflection+coloring+extension | cw. rotation | 13949 | ||||
| reflection+rotation+coloring | ccw. rotation | 13831 | ||||
| Test | translation+rotation+coloring | horizontal reflection | 26329 | |||
| translation+rotation+extension | vertical reflection | 26252 | ||||
Training on static data.
In addition to the model trained via MLC on a stream of dynamically changing visual interpretation grammars, as described in Section 3.2, we adopt the approach of Lake (2019) and train a transformer-based encoder-decoder on a dataset governed by a fixed visual grammar (referred to as basic seq2seq). Instead of episodes containing multiple study examples and queries, this dataset comprises individual input-output grid pairs, where the objective is to predict the output grid corresponding to a given input grid. This more closely resembles a standard training approach.
We construct a dataset of 1,300 grid pairs, partitioned into 1,260 training samples, 20 validation samples, and 20 test samples. All examples conform to a fixed visual interpretation grammar, meaning that a set of fixed visual indicators consistently maps to specific transformations. Samples represent primitive transformations, as well as level-1 and level-2 transformation compositions. As with our other experiments, the test set includes level-2 transformation compositions that were not observed during training—only their constituent components and level-1 compositions were seen during training. For instance, the test set might include transformations composed of shape-based downward translation, color-based horizontal reflection, and neighbor-based upward extension. However, only their decomposed elements have been shown during training.
The model is trained for 200 episodes using the parameters specified in Section B. While it successfully fits the training data (accuracy of over 99%), it fails to generalize to the test set, achieving a test accuracy of 0.0%.
For the prediction you need to understand the transformations displayed in the provided examples and apply them to the final input grid.
### Your Task:
1.
Analyze the example pairs to infer the transformation rules applied to each input grid.
2.
Identify how these transformations are applied to generate the output grids.
3.
Apply the deduced transformations to the final input grid.
4.
Output the correctly transformed 10x10 grid.
### Output Requirements:
Return only the final output grid.
Do not include any extra text, explanations, or comments.
The output must be formatted exactly as: ‘output: [[…]]‘
The output grid must be a 10x10 list of lists containing only integers between 0 and 9 (inclusive).
Do not include unnecessary line breaks or additional text beyond the specified format.
### Input Format:
You will receive the following data:
1.
Study examples: A list of 3 few-shot example pairs, formatted as:
‘example input 1: [[…]], example output 1: [[…]], …, example input 3: [[…]], example output 3: [[…]]‘
2.
Final input: A single 10x10 list of lists on which you must apply the inferred transformation(s).
Your goal is to determine the correct transformation and return the final output grid.
### Input:
Study examples:
example input 1: 2-dimensional array representing the input grid of example 1
example output 1: 2-dimensional array representing the output grid of example 1
example input 3: 2-dimensional array representing the input grid of example 3
example output 3: 2-dimensional array representing the output grid of example 3
Final input: 2-dimensional array representing the final query input grid
For the prediction you need to understand the transformations displayed in the provided examples and apply them to the final input grid.
### Your Task:
1.
Analyze the example pairs to infer the transformation rules applied to each input grid.
2.
Identify how these transformations are applied to generate the output grids.
3.
Apply the deduced transformations to the final input grid.
4.
Output the correctly transformed 10x10 grid.
### Output Requirements:
Return only the final output grid.
Do not include any extra text, explanations, or comments.
The output must be formatted exactly as: ‘output: [[…]]‘
The output grid must be a 10x10 list of lists containing only integers between 0 and 9 (inclusive).
Do not include unnecessary line breaks or additional text beyond the specified format.
### Input Format:
You will receive the following data:
1.
Study examples: A list of 3 few-shot example pairs, formatted as:
‘example input 1: [[…]], example output 1: [[…]], …, example input 3: [[…]], example output 3: [[…]]‘
2.
Final input: A single 10x10 list of lists on which you must apply the inferred transformation(s).
3.
Image input: Additionally, you receive an image that visualizes the 3 few-shot example pairs and the final input query.
Your goal is to determine the correct transformation and return the final output grid.
### Input:
Study examples:
example input 1: 2-dimensional array representing the input grid of example 1
example output 1: 2-dimensional array representing the output grid of example 1
example input 3: 2-dimensional array representing the input grid of example 3
example output 3: 2-dimensional array representing the output grid of example 3
Final input: 2-dimensional array representing the final query input grid
The first 6 example pairs demonstrate primitive transformations based on the object’s color, shape, or the presence of an additional object. For instance, objects of a certain color within the 10x10 input grid might undergo a translation, while objects of a certain shape (distinct numerical pattern) are being rotated.
The latter 6 example pairs involve composite transformations, meaning multiple transformations are applied simultaneously. For instance, for objects that have the appropriate color and shape, both a translation and rotation are applied simultaneously.
For the final prediction you need to understand and further combine the transformations displayed in the provided examples and apply them to the final input grid.
### Your Task:
1.
Analyze the example pairs to infer the transformation rules applied to each input grid.
2.
Identify how these transformations might combine to generate the output grids.
3.
Apply the deduced transformations to the final input grid.
4.
Output the correctly transformed 10x10 grid.
### Output Requirements:
Return only the final output grid.
Do not include any extra text, explanations, or comments.
The output must be formatted exactly as: ‘output: [[…]]‘
The output grid must be a 10x10 list of lists containing only integers between 0 and 9 (inclusive).
Do not include unnecessary line breaks or additional text beyond the specified format.
### Input Format:
You will receive the following data:
1.
Study examples: A list of 12 study example pairs, formatted as:
‘example input 1: [[…]], example output 1: [[…]], …, example input 12: [[…]], example output 12: [[…]]‘
2.
Final input: A single 10x10 list of lists on which you must apply the inferred transformation(s).
Your goal is to determine the correct transformation and return the final output grid.
### Input:
Study examples:
example input 1: 2-dimensional array representing the input grid of example 1
example output 1: 2-dimensional array representing the output grid of example 1
example input 12: 2-dimensional array representing the input grid of example 12
example output 12: 2-dimensional array representing the output grid of example 12
Final input: 2-dimensional array representing the final query input grid
The first 6 example pairs demonstrate primitive transformations based on the object’s color, shape, or the presence of an additional object. For instance, objects of a certain color within the 10x10 input grid might undergo a translation, while objects of a certain shape (distinct numerical pattern) are being rotated.
The latter 6 example pairs involve composite transformations, meaning multiple transformations are applied simultaneously. For instance, for objects that have the appropriate color and shape, both a translation and rotation are applied simultaneously.
For the final prediction you need to understand and further combine the transformations displayed in the provided examples and apply them to the final input grid.
### Your Task:
1.
Analyze the example pairs to infer the transformation rules applied to each input grid.
2.
Identify how these transformations might combine to generate the output grids.
3.
Apply the deduced transformations to the final input grid.
4.
Output the correctly transformed 10x10 grid.
### Output Requirements:
Return only the final output grid.
Do not include any extra text, explanations, or comments.
The output must be formatted exactly as: ‘output: [[…]]‘
The output grid must be a 10x10 list of lists containing only integers between 0 and 9 (inclusive).
Do not include unnecessary line breaks or additional text beyond the specified format.
### Input Format:
You will receive the following data:
1.
Study examples: A list of 12 study example pairs, formatted as:
‘example input 1: [[…]], example output 1: [[…]], …, example input 12: [[…]], example output 12: [[…]]‘
2.
Final input: A single 10x10 list of lists on which you must apply the inferred transformation(s).
3.
Image input: Additionally, you receive an image that visualizes the 12 study example pairs and the final input query.
Your goal is to determine the correct transformation and return the final output grid.
### Input:
Study examples:
example input 1: 2-dimensional array representing the input grid of example 1
example output 1: 2-dimensional array representing the output grid of example 1
example input 12: 2-dimensional array representing the input grid of example 12
example output 12: 2-dimensional array representing the output grid of example 12
Final input: 2-dimensional array representing the final query input grid